OpenStudy (anonymous):

When power cables are suspended between two towers they form a curve called the catenary shown below. The equation of the curve is given by: y=acosh(x/a) where the distance between the two towers is 2b. Find the slope of the catenary where the cable meets the right hand tower.

OpenStudy (anonymous):



OpenStudy (anonymous):

that is the drawing that it gives me

OpenStudy (anonymous):

I am not sure how to start it

OpenStudy (inkyvoyd):

d/dx sinh(x)=cosh(x) d/dx cosh(x)=-sinh(x) to derive these formulas use the equivalent cosh(x)=(e^x+e^-x)/2 and sinh(x)=(e^x-e^-x)/2

OpenStudy (anonymous):

\[y=a(\frac{ e^(x/a)+e^(-x/a) }{ 2 })\]

OpenStudy (anonymous):

the (x/a)'s are exponents

OpenStudy (anonymous):

does x=b? since I am trying to find the slope of the right side

OpenStudy (inkyvoyd):

yeah prolly

OpenStudy (inkyvoyd):

but there's no need to use the expanded form. just do d/dx cosh(x)=-sinh(x)